Razor:

I got another math assignment for people to answer muhahaha e,e mainly dude tho if you dont mind

Razor:

1 attachment
Razor:

@dude

Razor:

:/

dude:

Was busy my bad Anyway to find the new measurements multiply the measurement given by the scale factor

Razor:

15 + 8 x 1/2 or am I wrong lmao @dude

Razor:

I'm actually multi-tasking with artwork and this right now so I'm not gonna reply as fast as I usually do

jhonyy9:

ecery lengths multiplie by this scale factor 15×1/2 = ? this will be the new length and 8×1/2 = ? this will be the new width

jhonyy9:

ecery wan being every sorry

jhonyy9:

do you understand it ?

jhonyy9:

hope you know when you multiplie suppose 10 by 1/2 so this mean 10 halfed what result 5 hope this example help you

Razor:

XD Oh ok so 1/2 is supposed to be multiplied by the width and the length

jhonyy9:

yes exactly

jhonyy9:

hope helped bye

Razor:

bye ^^

jhonyy9:

or do you have questions ?

Razor:

well I thought you had to go lmao wait ok so 1/2 multiplied by both of those is 7.5 for 15 and 4 for 8

jhonyy9:

yes right

jhonyy9:

the new rectangle will have length of 7,5 and width of 4 hope now is clearly sure

Razor:

So That would kind of be the answer for A) in my image that I posted??

jhonyy9:

you need drawing a new rectangle with AB and CD = 7,5 and with BD and AC = 4 ok. ?

Razor:

ok sorry for not responding I'm working on multiple things at once

jhonyy9:

np

Razor:

Okay how do I make a rectangle with AB and CD = to 7.5 and BD and AC = to 4.. I already drew the rectangle on a scrap piece of paper soo

Razor:

Um hello?? @jhonyy9

Shadow:

Yeah that's all you have to do (a) \[15 \times \frac{ 1 }{ 2 } = 7.5 cm\] \[8 \times \frac{ 1 }{ 2 } = 4 cm\] (b) 

Shadow:

To fulfill your requirements, switch the positions of D' and C' in my case.

Razor:

lmao ok thanks @Shadow

Shadow:

No problem